OpenStudy (anonymous):

Erica bought 3 1/2 yards of fabric. If she uses 2/3 of the fabric to make a curtain, how much will she have left?

OpenStudy (darthvader2900):

3 1/2 - (2/3)(3 1/2) = 7/2 - (2/3)(7/2) = 7/2 - 7/3 = 21/6 - 14/6 = 7/6 or 1 1/6 yds

OpenStudy (darthvader2900):

ok?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{erica brougth }3\cfrac{1}{2} \normalsize \text{yards of fabric } \hspace{.33em}\\~\\ \normalsize \text{she uses }\cfrac{2}{3} \normalsize \text{of the fabric to make a curtain } \hspace{.33em}\\~\\ \normalsize \text{so remaining fabric }=3\cfrac{1}{2}-\cfrac{2}{3}3\cfrac{1}{2} \hspace{.33em}\\~\\ =\cfrac{3\times 2+1}{2}-\cfrac{2}{3}\times \cfrac{3\times 2+1}{2} \hspace{.33em}\\~\\ =\cfrac{7}{2}-\cfrac{2}{3}\times \cfrac{7}{2} \hspace{.33em}\\~\\ =\cfrac{7}{2}\left(1-\cfrac{2}{3}\right) \hspace{.33em}\\~\\ =\cfrac{7}{2}\left(\cfrac{3-2}{3}\right) \hspace{.33em}\\~\\ =\cfrac{7}{2}\left(\cfrac{1}{3}\right) \hspace{.33em}\\~\\ =\cfrac{7}{6} \hspace{.33em}\\~\\ \end{align}}\)